Obtain a relation for the distance travelled by an object moving with a uniform acceleration in the interval between 4th and 5th seconds.
1
Expert's answer
2017-07-28T15:59:06-0400
s_5=5u+(a5^2)/2=5u+12.5a
s_4=4u+(a4^2)/2=4u+8a
Δs=(5u+12.5a)-(4u+8a)=u+4.5a.